OpenStudy (anonymous):

Factor 6x²+x-2. Explain your steps!

OpenStudy (anonymous):

(3x+2)(2x-1) you first need to know the factors of the terms in the equation. Once you know these, you're half way there 3x*2x=6x^2 3s*(-1)=-3x 2*2x=4x 2*(-1)=-2 there is 4x and -3x which gives you x notice you have 6x^2+x-2 ... so your answer is correct. when multiplying... you multiply the first term with the first time of the other brackets and then, then with the second term in that bracket then u multiply the 2nd term in first brack with 1st term in second bracket, then with 2nd term makes sense?

OpenStudy (anonymous):

Yess, omg thank you! :)